OpenStudy (rianiscool):

need help asap in math Identify the property shown. 6j + (2k + 3) = (6j + 2k) + 3 A. associative B. commutative C. distributive D. transitive

OpenStudy (igreen):

Associative property examples: \(a + (b + c) = (a + b) + c\) \(a(bc) = (ab)c\)

OpenStudy (igreen):

Therefore, Associative property is the correct answer.

OpenStudy (rianiscool):

are you shure